OpenStudy (anonymous):

Traveling at 55 miles per hour, how many minutes, rounded to the nearest whole number, does it take to drive 310 miles from Boston to Syracuse? A(6 B(338 C(365 D(3,300

OpenStudy (anonymous):

1 hr = 60 minutes

OpenStudy (anonymous):

55/60 = 310/x --- 55 miles to 60 minutes = 310 miles to x minutes

OpenStudy (anonymous):

(55)(x) = (60)(310) 55x = 18600 x = 18600/55 x = 338 2/11 to the nearest whole number is 338

OpenStudy (andrearchat):

\[55\div60=x\] then do \[x \times310\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Np (: Don't forget to medal or fan